OpenStudy (anonymous):

The length of the bridge is 48 ft and the height is 15 ft. It is made of 5 congruent triangles. Are these triangle acute, right, or obtuse? How many feet of material is needed to build one side of the bridge?

OpenStudy (anonymous):

this is the diagram

OpenStudy (anonymous):

please help me with this! it is really important

OpenStudy (anonymous):

The triangles are acute.

OpenStudy (anonymous):

sstarica might be showing why...

OpenStudy (anonymous):

lol, I was about to say that they are all the same that's all

OpenStudy (anonymous):

and was about to ask you why LOL

OpenStudy (anonymous):

oh :)

OpenStudy (anonymous):

^_^ proceed.

OpenStudy (anonymous):

Okay QEvelyn, triangles are acute when each of their internal angles are less the 90 degrees. So the challenge here is to determine the angles. Here we go...

OpenStudy (anonymous):

You have three triangles on the bottom, and the sum of their bases equals 48. The base of each triangle is then, 48/3 = 16

OpenStudy (anonymous):

You're also told the height, which is 15. This is the same height you'd find from the base of the triangle to the apex. If you draw a perpendicular line from the apex (tip) to the bottom, it will bisect the base. You'll then have, in each triangle, two right-angled triangles with base = 8 height = 15 The hypotenuse is the side of the bigger triangle it came from. You need to find that. You can use Pythagoras' Theorem for that. So, \[s^2=8^2+15^2=289 \rightarrow s=17\]

OpenStudy (anonymous):

lol myininaya, you've answered in the wrong place

OpenStudy (anonymous):

oh nvm >_< her answer is gone

OpenStudy (anonymous):

Since two of these triangles make up your larger one, because of symmetry, you have an isosceles triangle (putting the two right-angled ones back together again). The base is 16 and the sides are 17 each. Because the triangle is isosceles, the two angles at the base will be equal. Call them \[\alpha\]Let's call the angle at the top\[\theta\]By the law of cosines, \[16^2=17^2+17^2-2 \times 17^2 \cos \theta \rightarrow \cos \theta = \frac{161}{289}\]which implies\[\theta \approx 56.145^o\]Since the sum of the angles of a plane triangle equals 180 degrees, and since the triangle is isosceles, you have\[180^o=2 \alpha +56.145^o \rightarrow \alpha \approx 61.928^o\]Your three angles in each of the bottom triangles is then\[56.145^o, 61.928^o,61.928^o\]All of these angles are less than 90 degrees, so the triangle is acute.

OpenStudy (anonymous):

Since you're given that the triangles are congruent, proving this for one proves for all.

OpenStudy (anonymous):

and he strikes again ~ :)

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

QEvelyn, clear? :)

OpenStudy (anonymous):

I think the questioner has fled.

OpenStudy (anonymous):

was afraid of all the writing lol

OpenStudy (anonymous):

you've scared me, but when I read it, it was all clear :)

OpenStudy (anonymous):

scared you?

OpenStudy (anonymous):

the HUGE reply came out of nowhere in a BLINK ._. LOL

OpenStudy (anonymous):

Problem is, all this writing can be done away with if you could draw on this thing.

OpenStudy (anonymous):

Haha

OpenStudy (anonymous):

yeah , I know lol ^_^

OpenStudy (anonymous):

he's prolly checking your reply now :)

OpenStudy (anonymous):

Do you want to finish it - I want to go to bed?!

OpenStudy (anonymous):

but it's already done?!

OpenStudy (anonymous):

There's something about area.

OpenStudy (anonymous):

well add up all the areas then,right?

OpenStudy (anonymous):

A = 120x5 = 600 ft^2 done!

OpenStudy (anonymous):

he'll need that much for that, I guess

OpenStudy (anonymous):

How many feet of material or something...

OpenStudy (anonymous):

wait, it's not .

OpenStudy (anonymous):

Could be one of those frame bridges.

OpenStudy (anonymous):

I figured the answer till the area of each, but not the material

OpenStudy (anonymous):

I get 182 feet for one side.

OpenStudy (anonymous):

Just check, I'm fried.

OpenStudy (anonymous):

lol, alright scoot

OpenStudy (anonymous):

Bye! Happy mathing.

OpenStudy (anonymous):

lol bye, I'll try to solve it though ^_^

OpenStudy (anonymous):

I can't seem to figure out how he got 182 ft

OpenStudy (anonymous):

sstarica, you'll kick yourself: it's just summing up each of the lengths.

OpenStudy (anonymous):

HI GUYS thank you very much it went all over my head and sorry im so late to respond. the teacher will go over it in class next week and ill tell you all what she got. thank you again and i know who to come to again for math help! oh yeah and im a she :)

OpenStudy (anonymous):

And I just became a fan of both of you so Iokisan, you are now a Hero!!!

OpenStudy (anonymous):

Oh, so you're the one who clicked me over! Thank you!